A box contains 25% black balls and 75% white balls. The same number of black balls as were in the box are added (so new number of black balls is twice the original number). A ball is now drawn from the box at random. What is the probability that it is black?
Answer by stanbon(75887)

A box contains 25% black balls and 75% white balls. The same number of black balls as were in the box are added (so new number of black balls is twice the original number). A ball is now drawn from the box at random. What is the probability that it is black?
-----
Converting the problem to numbers:
Original mix: 25 black and 75 white.
--
New mix: 50 black and 75 white
---
P(black) = 50/125 = 2/5 = 40%
====================================
Cheers,
Stan H.